# MENTOR: Efficient Multimodal-Conditioned Tuning for Autoregressive Vision Generation Models

### v1

@@ 3.1 Experimental Setup
Implementation Details The multimodal encoder is initialized using CLIP-Large-Patch14 (Radford et al., 2021) and Flant5-XL (Chung et al., 2022), with a 224×224 image receptive field. The generator is initialized from LlamaGen-XL (Sun et al., 2024b)(775M). A two-layer MLP serves as the projector. We freeze the encoder, training only the projector and generator for one epoch in Stage 1, then fine-tune the full model (except the vision encoder) for two more epochs in Stage 2. Training on 8 A100 GPUs (80 GB each) takes about 1.5 days. More details are in Appendix B.
Benchmark & Metric. We evaluate Mentor on DreamBench (Ruiz et al., 2023) and DreamBench++ (Peng et al., 2025) benchmarks. DreamBench employs CLIP and DINO scores to measure the fidelity to images and prompts. DreamBench++ offers a scaled and diverse evaluation dataset and introduces a human-aligned automatic evaluation protocol using GPT-4o, addressing limitations of DreamBench evaluation. GPT-4o evaluator scores generations on two axes: Concept Preservation (CP), measuring the retention of the subject’s visual identity, and Prompt Following (PF), evaluating how accurately the image reflects the text prompt. Details can be found in the § E.1.
Baselines. We compare our method against various baselines, categorized as follows:
  • Fine-tuning-based methods: Textual Inversion (Gal et al., 2022), and DreamBooth (Peng et al., 2025), which fine-tune models with auxiliary control mechanisms on a subset of benchmark data.
  • Test Time Tuning-Free Methods: Models pretrained on large-scale data and evaluated in a zero-shot. Diffusion-based models like BLIP-Diffusion (Li et al., 2023a), Emu2 (Sun et al., 2024c), variants of IP-Adapter (Ye et al., 2023), and DreamEngine (Chen et al., 2025), as well as autoregressive models like Unified-IO 2 (Lu et al., 2023) and Lumina-mGPT (Zhuo et al., 2024).
3.2 Main results
Table 3 comprehensively evaluates our proposed autoregressive (AR) framework on the DreamBench++ benchmark, comparing it with diffusion-based and autoregressive-based baselines. Mentor demonstrates highly competitive performance, particularly in achieving a strong balance between the guidance of both input modalities. Notably, this is achieved despite utilizing significantly fewer training resources and suboptimal model components compared to the state-of-the-art baselines.
@@ 3.2 Main results
[…]
@@
Textual Inv. SD v1.5 - 860M 0.50 0.36 0.31 0.36 0.38 0.67 0.69 0.44 0.62 0.24 0.61
DreamBooth SD v1.5 - 860M 0.64 0.20 0.49 0.48 0.49 0.79 0.78 0.50 0.72 0.36 0.68
DreamBooth-L SDXL v1.0 - 2.60B 0.75 0.31 0.54 0.72 0.60 0.90 0.90 0.75 0.87 0.52 0.69
@@
Unified-IO2* Unified-IO2 8.5B 7.00B 0.77 0.80 0.64 0.82 0.72 0.24 0.18 0.11 0.19 0.14 3.79
Lumina-mGPT Chameleon 10M 7.00B 0.95 0.97 0.89 0.85 0.91 0.31 0.25 0.15 0.25 0.23 3.64
@@
BLIP-Diffusion SD v1.5 130M 1.56B 0.67 0.56 0.47 0.51 0.55 0.58 0.51 0.30 0.50 0.27 1.10
Kosmos-G SD v1.5 200M 3.00B 0.62 0.63 0.46 0.57 0.54 0.48 0.62 0.41 0.51 0.28 1.06
IP-A-Plus ViT-H SDXL v1.0 10M 3.00B 0.90 0.85 0.76 0.91 0.83 0.50 0.38 0.28 0.41 0.34 2.02
Emu2 SDXL v1.0 16M 37.00B 0.67 0.55 0.45 0.45 0.53 0.73 0.72 0.56 0.69 0.36 0.77
@@
Mentor LlamaGen 3M 2.31B 0.65 0.36 0.57 0.47 0.55 0.86 0.85 0.80 0.84 0.47 0.65
Training Efficiency. A notable advantage of Mentor lies in its training efficiency. It is trained on only 3 million image-text pairs across two stages, substantially less than leading baselines, such as Emu2 (16M), Kosmos-G (200M), and DreamEngine (21M). Beyond the reduced data requirements, the training process is highly resource-efficient: the entire training process completes in 1.5 days with 8 GPUs. This contrasts sharply with other baselines, such as Kosmos-G, which necessitates 256 GPUs over three days. Despite this dramatically reduced computational and data budgets, Mentor achieves SOTA performance with balanced performance, highlighting its efficiency and effectiveness. Furthermore, Mentor remains highly competitive in size compared to larger counterparts like Emu2 (37B parameters) and DreamEngine (10.5B parameters), highlighting our framework’s effectiveness.
@@
3.3 Ablation Study
@@
Importance of Stage 1: Foundational Multimodal Alignment. As shown in Table 4, removing Stage 1 leads to the most severe performance drop, underscoring its foundational role. On DreamBench++, the CP score drops from 0.555 to 0.179, indicating a major loss in visual identity preservation, with PF also significantly reduced. Similar trends appear on DreamBench. Ablating only object segmentation task in Stage 1 (w/o Obj. Seg. in Stage 1) also hampers model performance. While the remaining image reconstruction task supports pixel-level alignment, allowing for reconstruction of input images, it inadvertently leads the model to exhibit a copy-paste behavior, failing to capture semantic and visual information of input images. It confirms that reconstruction alone is insufficient for robust multimodal alignment. Overall, these results highlight Stage 1 as critical for aligning multimodal inputs with output images. Without such alignment, the model struggles to ground visual concepts from images, severely impairing its visual preservation ability.
Contributions of Different Training tasks in Stage 2. The distinct contributions of different Stage 2 tasks are also evident in Table 4. Excluding the Image Recovery task leads to a sharp imbalance: while visual preservation metrics (CP, DINOv1, CLIP-I) show a notable increase, instruction following ability (PF and CLIP-T score) critically drops, showing an overfitting to visual features. This underscores that image recovery acts as a critical regularizer, encouraging the model to reconstruct incomplete visual contexts guided by text prompt, thereby fostering a balanced use of different modalities. Conversely, ablating either the Object Segmentation or the Subject-Driven Image generation significantly degrades visual preservation ability, as these tasks prompt the model to utilize the visual features of the input image effectively to generate images. Without these tasks, the model tends to over-rely on textual prompts, resulting in a regression towards a standard T2I model that neglects the reference image. These results highlight the importance of all Stage 2 tasks: image recovery ensures cross-modal balance, while object segmentation and subject-driven generation enhance the model’s ability to extract and utilize detailed visual information for image generation.
@@
3.4 Analysis
Efficiency and Effectiveness: AR vs. Diffusion. To evaluate the efficiency of our AR framework against diffusion-based approaches, we conducted a controlled comparison with Kosmos-G (Pan et al., 2024), a representative LMM-augmented diffusion model.Both models were trained from similar initializations on the same training data to ensure a fair comparison. Despite Kosmos-G employing a superior SD1.5 generator and Kosmos-1 as encoder, Mentor, which utilizes a underperformed LlamaGen generator and a FlanT5 based encoder, demonstrated significantly better performance on DreamBench++ as shown in Table 3. It highlights the effectiveness of Mentor in fostering efficient multimodal learning and strong multimodal conditional generation.
Comparison of Proposed Architecture Variants. As shown in Table 4, we compare two architectural variants of Mentor: the MLP-based connector (Mentor) and a Query-based connector for the multimodal encoder. It reveals that an MLP-based connector significantly outperforms the query-based variant in CP scores, especially for humans and objects. It suggests that after visual token compression, the query-based approach struggles to retain fine-grained visual details, which are crucial for generative fidelity, even when guided by textual queries. Nevertheless, due to effective token distillation, the Query-based variant facilitates training with multiple contextual images with minimal computational resources. Despite these differences, both variants exhibit competitive performance compared to other baselines in Table 3. This demonstrates that our simple and coherent architecture can effectively propagate visual features, whether or not perform aggressive visual token compression, highlighting the flexibility and robustness of our framework.
@@ 3.4 Analysis
Animal Human Object Style Overall Photorealistic Style Transfer Imaginative Overall
@@
Table 4: Ablation studies on architecture design and multi-image training
Method DreamBench++ DreamBench
@@
w. Multi-image $0.586\pm 0.006$ $0.829\pm 0.005$ $\mathbf{0.486}$ $72.487\pm 0.147$ $81.857\pm 0.152$ $28.545\pm 0.043$
[…]
@@
Image Reconstruction Fidelity. To quantify visual detail preservation in our framework, we evaluate Mentor on the Image Reconstruction Benchmark (Chen et al., 2025), which measures similarity between input and reconstructed images. After fine-tuning on reconstruction task for 1,000 steps, we compare the generated outputs with their originals using pixel-space $\ell_{2}$ distance, following pervious work (Chen et al., 2025). As shown in Table 5, Mentor outperforms strong baselines with comparable architectures—SeedTokenizer (Ge et al., 2023), EMU2 (Sun et al., 2024c), SeedX (Ge et al., 2024), and DreamEngine (Chen et al., 2025)—all of which couple LMMs with diffusion backbones. Mentor achieves the best reconstruction quality, exceeding the second-best by 50%, even with a $224{\times}224$ receptive field, while others varied from 384x384 to 512x512. These gains confirm our model’s effectiveness at conditioning on—and faithfully reproducing—visual inputs.
Versatility Across Different Multimodal Tasks. To explore broader applicabilities of our framework, we evaluate its adaptability across diverse generation tasks, including image segmentation, multi-image generation and multimodal in-context image generation. This was achieved with brief fine-tuning on relevant datasets, as detailed in Appendix G. Qualitative results in Figure 1 show that the Mentor produces coherent, high-quality outputs that adhere to the provided constraints without requiring any architectural modifications. While achieving performance in each specific domain would necessitate more specialized training and potentially more powerful multimodal encoder and generator components, these initial results underscore our framework’s versatility and its potential as an effective foundation for a variety of multimodal conditional image generation applications.
@@ 4.1 Image Generation with Complex Multimodal Control
Researchers have developed image generation via diffusion models conditioned on multimodal inputs like canny edges (Zhang et al., 2023) and reference images (Zhao et al., 2024; Meng et al., 2022). ControlNet (Zhang et al., 2023) uses auxiliary parameters, while Mix-of-Show (Gu et al., 2023) and FLUXSynID (Ismayilov et al., 2025) use LoRA modules for multi-concept control and identity preservation. DreamBooth (Ruiz et al., 2023) enable subject-specific fine-tuning but limit generalization. SuTI (Chen et al., 2023b) address it with scalable data and training. To enhance flexibility, recent work integrates LMMs with diffusion models by mapping LMM embedding into diffusion spaces (Koh et al., 2023; Sun et al., 2023; Dong et al., 2023; Li et al., 2024). Approaches like Kosmos-G (Pan et al., 2024), Emu-2 (Sun et al., 2024c), Seed-X (Ge et al., 2024), and DreamEngine (Chen et al., 2025) explore more complex multimodal prompt and fine-grained multimodal control. Yet, balancing guidance from diverse modalities remains a core challenge (Han et al., 2024; Ye et al., 2023; Mao et al., 2024). EMMA (Han et al., 2024) employs a gated perceiver resampler for dynamic signal integration, while RealCustom++(Mao et al., 2024) disentangles subject identity and textual fidelity via cross-layer projectors. OmniControl (Tan et al., 2025) introduces a bias term into multimodal attention. Nonetheless, these method often require substantial computational resources, and achieving efficient, robust, and scalable multimodal integration remains an open problem.
4.2 Autoregressive Multimodal Image Generation
@@
Autoregressive models have driven progress in T2I generation, from DALL·E Ramesh et al. (2021) and Parti (Yu et al., 2022) to LlamaGen (Sun et al., 2024b) and GPT4O (OpenAI, 2024). Recent work extends it to multimodal settings: Models like Chameleon (Team, 2024b), LWM Liu et al. (2024b), AnyGPT (Zhan et al., 2024), and EMU3 (Wang et al., 2024b) treat text and images as unified token sequences via early-fusion transformers, yet still emphasize text-to-image generation with limited support for multimodal conditioning. Janus (Wu et al., 2024a) decouples visual understanding and generation via distinct pathways, lacks support for multimodal image generation. MUSE-VL Xie et al. (2025) and VILA-U (Wu et al., 2024c) align discrete visual tokens with text to improve perception, but remain oriented toward understanding tasks rather than image generation. Unified-IO2 (Lu et al., 2023) is trained autoregressively from scratch for both understanding and generation across modalities, while Lumina-mGPT Zhuo et al. (2024) enhances Chameleon with omnipotent supervised fine-tuning for broader multimodal tasks. Nonetheless, these models often over-rely on visual inputs while ignoring textual prompts. Overall, while models like VILA-U Wu et al. (2024c), EMU3 Wang et al. (2024b), and Janus Wu et al. (2024a) have advanced text-to-image generation, robust multimodal conditional image generation remains an open and underexplored challenge.

### v2

@@ 3.1 Experimental Setup
Method T2I Model Train Data Model Size Concept Preservation (CP) Prompt Following (PF) $\cdotpPF CP$/PF
@@
Textual Inv. SD v1.5 - 860M 0.50 0.35 0.30 0.36 0.38 0.68 0.70 0.44 0.63 0.24 0.60
DreamBooth SD v1.5 - 860M 0.64 0.19 0.49 0.48 0.49 0.79 0.78 0.51 0.72 0.36 0.68
DreamBooth-L SDXL v1.0 - 2.60B 0.75 0.31 0.54 0.71 0.60 0.90 0.90 0.75 0.87 0.52 0.69
@@
VARGPT-v1.1 Infinity (VAR) 8.3M 9.00B 0.25 0.13 0.14 0.27 0.19 0.31 0.48 0.33 0.37 0.07 0.51
Unified-IO2* Unified-IO2 8.5B 7.00B 0.77 0.80 0.64 0.82 0.72 0.24 0.18 0.11 0.19 0.14 3.74
Lumina-mGPT Chameleon 10M 7.00B 0.95 0.97 0.89 0.85 0.91 0.31 0.25 0.15 0.25 0.23 3.63
OmniGen* OmniGen 700M 3.80B 0.39 0.35 0.25 0.22 0.30 0.70 0.71 0.67 0.70 0.21 0.43
@@
BLIP-Diffusion SD v1.5 130M 1.56B 0.67 0.56 0.47 0.51 0.55 0.58 0.52 0.30 0.50 0.27 1.11
Kosmos-G SD v1.5 200M 3.00B 0.62 0.64 0.46 0.56 0.55 0.48 0.62 0.39 0.51 0.28 1.07
IP-A-Plus ViT-H SDXL v1.0 10M 3.00B 0.90 0.84 0.76 0.91 0.84 0.50 0.39 0.28 0.41 0.35 2.02
Emu2 SDXL v1.0 16M 37.00B 0.67 0.55 0.45 0.44 0.53 0.73 0.73 0.56 0.69 0.36 0.76
@@
Mentor LlamaGen 3M 2.31B 0.65 0.36 0.57 0.47 0.56 0.86 0.85 0.80 0.84 0.47 0.66
Implementation Details We initialized the multimodal encoder with CLIP-Large-Patch14 (Radford et al., 2021) and FlanT5-XL (Chung et al., 2022), which has 224×224 image receptive field, and the generator with LlamaGen-XL (Sun et al., 2024a). Training on 8 A100 GPUs about 1.5 days. More Details in Appendix˜B and §˜I.2.
Benchmark & Metric. We evaluate Mentor on DreamBench (Ruiz et al., 2023) and DreamBench++ (Peng et al., 2025) benchmarks. DreamBench uses CLIP and DINO scores to assess image fidelity and prompt alignment. DreamBench++ addresses limitations of DreamBench evaluation and evaluate on two axes: Concept Preservation (CP), measuring the retention of the subject’s visual identity, and Prompt Following (PF), evaluating how accurately the image reflects the text prompt. A human evaluation study confirms high consistency between GPT-based metrics and human judgments. Further details are provided in §˜E.1.
Baselines. We compare Mentor with both fine-tuning and tuning-free methods. Fine-tuning baselines include Textual Inversion (Gal et al., 2022) and DreamBooth (Peng et al., 2025). Tuning-free baselines include diffusion-based models—BLIP-Diffusion (Li et al., 2023a), Emu2 (Sun et al., 2024c), IP-Adapter (Ye et al., 2023), OmniGen (Xiao et al., 2024b), and DreamEngine (Chen et al., 2025b)—and AR models such as Unified-IO 2 (Lu et al., 2023), Lumina-mGPT (Zhuo et al., 2024), and VARGPT-v1.1 (Zhuang et al., 2025).
3.2 Main Results
Table˜1 comprehensively evaluates our proposed autoregressive (AR) framework on the DreamBench++ benchmark, comparing it with diffusion-based and autoregressive-based baselines. Mentor demonstrates highly competitive performance, particularly in achieving a strong balance between the guidance of both input modalities. Notably, this is achieved despite utilizing significantly fewer training resources and suboptimal model components compared to the state-of-the-art baselines.
@@ 3.2 Main Results
Training Efficiency. A notable advantage of Mentor lies in its training efficiency. It is trained on only 3 million image-text pairs across two stages, substantially less than leading baselines, such as Emu2 (16M), Kosmos-G (200M), and DreamEngine (21M). Beyond the reduced data requirements, the training process is highly resource-efficient: the entire training process completes in 1.5 days with 8 GPUs. This contrasts sharply with other baselines, such as Kosmos-G, which necessitates 256 GPUs over three days. Despite this dramatically reduced computational and data budgets, Mentor achieves competitive performance with balanced performance, highlighting its efficiency and effectiveness. Furthermore, Mentor remains highly competitive in size compared to larger counterparts, highlighting our effectiveness.
@@
3.3 Ablation Study
[…]
Importance of Stage 1: Foundational Multimodal Alignment. As shown in Table˜2, removing Stage 1 leads to severe performance drop, underscoring its foundational role. On DreamBench++, the CP score drops from 0.555 to 0.179, indicating a major loss in visual identity preservation, with PF also significantly reduced. Similar trends appear on DreamBench. Ablating only object segmentation task in Stage 1 (w/o Obj. Seg. in Stage 1) also hampers model performance. While the remaining image reconstruction task supports pixel-level alignment, allowing for reconstruction of input images, it inadvertently leads the model to exhibit a copy-paste behavior, failing to capture semantic and visual information of input images. It confirms that reconstruction alone is insufficient for robust multimodal alignment. Overall, these results highlight that without Stage 1, the model struggles to ground visual concepts from images, severely impairing its visual preservation ability.
@@ 3.3 Ablation Study
Animal Human Object Style Overall Photo. Style. Imag. Overall
@@
3.4 Analysis
Efficiency and Effectiveness: AR vs. Diffusion. To evaluate the efficiency of our AR framework against diffusion-based methods, we conducted a controlled comparison with Kosmos-G (Pan et al., 2024), a representative LMM-augmented diffusion model.Both models were trained from similar initializations on same training data to ensure a fair comparison. Shown in Table˜3, despite Kosmos-G employing a superior SD1.5 generator and Kosmos-1 encoder, Mentor, still shown significantly better performance with limited training data, highlighting the effectiveness of our framework.
@@ 3.4 Analysis
Comparison of Architecture Variants. We compare architectural variants of Mentor that differ in how visual features are connected to the generator. Since the multimodal encoder produces hundreds of visual tokens per image, the inflated context length poses significant computational challenges, particularly in multi-image settings. Token compression alleviates this issue but may compromise fine-grained visual fidelity. To navigate this trade-off, as shown in Table˜2, we evaluate a query-based connector that leverages a Query-Former Li et al. (2023b) to compress long visual sequences into fixed-size representations, guided by textual queries (Li et al., 2023a) highlighting important concepts in images. The results reveals that while the query-based approach substantially reduces computational cost, it struggles to preserve fine-grained visual details crucial for generative fidelity, even guided by textual queries. Consequently, the MLP-based connector(Mentor) achieves higher fidelity, especially for human and object regions, whereas the query-based variant offers a favorable trade-off between efficiency and performance. Despite these differences, both variants exhibit competitive performance compared to other baselines.
Table 5: Analysis on architecture design, multi-image training and reinforcement learning
Method DreamBench++ DreamBench
[…]
w. Multi-image $0.586\pm 0.006$ $0.829\pm 0.005$ $0.486$ $72.487\pm 0.147$ $81.857\pm 0.152$ $28.545\pm 0.043$
w. GRPO $0.609\pm 0.006$ $0.866\pm 0.004$ $\mathbf{0.527}$ $77.176\pm 0.199$ $83.343\pm 0.287$ $29.810\pm 0.059$
[…]
@@
Image Reconstruction Fidelity. To quantify visual detail preservation in our framework, we evaluate Mentor on the Image Reconstruction Benchmark (Chen et al., 2025b), which measures similarity between input and reconstructed images. After fine‑tuning on reconstruction task for 1,000 steps, we compare the generated outputs with their originals, following previous work (Chen et al., 2025b). As shown in Table˜4, Mentor outperforms strong baselines with comparable architectures—SeedTokenizer (Ge et al., 2023), EMU2 (Sun et al., 2024c), SeedX (Ge et al., 2024), and DreamEngine (Chen et al., 2025b)—all of which couple LMMs with diffusion backbones. Mentor achieves best reconstruction quality even with a $224{\times}224$ receptive field, while others varied from 384x384 to 512x512.
Versatility Across Different Multimodal Tasks. To explore broader applicabilities of our framework, we evaluate its adaptability across diverse tasks, including image segmentation, multi-image generation and multimodal in-context image generation. This was achieved with brief fine-tuning on relevant datasets, as detailed in Appendix˜G. Qualitative results in Figure˜2, Figure˜11, Figure˜14 and Figure˜15 show that Mentor produces coherent outputs that adhere to provided constraints without any architectural modifications. While achieving performance in each specific domain would necessitate more specialized training and potentially more powerful components, these initial results underscore our framework’s versatility and its potential as an effective foundation for multimodal image generation.
[…]
@@ 4.1 Image Generation with Complex Multimodal Control
Recent advances in diffusion models enable image generation conditioned on multimodal inputs such as canny edges (Zhang et al., 2023) or reference images (Zhao et al., 2024; Meng et al., 2022). DreamBooth (Ruiz et al., 2023) enables subject-specific fine-tuning but limits generalization. SuTI (Chen et al., 2023b) addresses it with scalable data and training. To enhance flexibility, recent work integrates LMMs with diffusion models (Koh et al., 2023; Sun et al., 2023; Dong et al., 2023; Li et al., 2024). Approaches like Kosmos-G (Pan et al., 2024), Emu-2 (Sun et al., 2024c), Seed-X (Ge et al., 2024), and DreamEngine (Chen et al., 2025b) explore more complex multimodal prompt and fine-grained multimodal control. Yet, balancing guidance from diverse modalities remains a core challenge (Han et al., 2024b; Ye et al., 2023; Mao et al., 2024). EMMA (Han et al., 2024b) employs a gated perceiver resampler for dynamic signal integration, while RealCustom++(Mao et al., 2024) disentangles subject identity and textual fidelity via cross-layer projectors. OmniControl (Tan et al., 2025) introduces a bias term into multimodal attention. Recently, large-scale systems such as Qwen-Image (Wu et al., 2025a), Begal (Deng et al., 2025), and Nano-Banana (Comanici et al., 2025) further demonstrate impressive multimodal generation capabilities through large-scale training on extensive multimodal corpora across thousands of GPUs. Nonetheless, these methods still demand substantial resources, and achieving efficient multimodal integration remains an open problem.
4.2 Autoregressive Multimodal Generation
@@
Autoregressive models have driven progress in T2I generation, from DALL·E Ramesh et al. (2021) to LlamaGen (Sun et al., 2024a) and GPT4O (OpenAI, 2024). Recent work extends it to multimodal settings: Models like Chameleon (Team, 2024a), LWM Liu et al. (2024a), AnyGPT (Zhan et al., 2024), and EMU3 (Wang et al., 2024b) treat text and images as unified token sequences via early-fusion transformers, yet still emphasize text-to-image generation with limited support for multimodal conditioning. Janus (Wu et al., 2024a) decouples visual understanding and generation via distinct pathways but lacks support for multimodal image generation. MUSE-VL Xie et al. (2025) and VILA-U (Wu et al., 2024c) align discrete visual tokens with text to improve perception, but remain oriented toward understanding tasks rather than image generation. Unified-IO2 (Lu et al., 2023) is trained autoregressively from scratch for both understanding and generation across modalities, while Lumina-mGPT Zhuo et al. (2024) and VARGPT (Zhuang et al., 2025) enhances Chameleon and VAR (Han et al., 2024a) with supervised fine-tuning for broader multimodal tasks. Overall, while models like VILA-U Wu et al. (2024c), EMU3 Wang et al. (2024b), and Janus Wu et al. (2024a) have advanced text-to-image generation, robust multimodal conditional image generation remains an open and underexplored challenge.